# Nonconvex Game and Multi Agent Reinforcement Learning for Zonal Ancillary Markets

### v2

@@ I-A Related Work
$\Delta_{n}^{min},\Delta_{n}^{max}$ Lower and upper capacity for producer $n$
@@
$\pi_{n}^{min},\pi_{n}^{max}$ Lower and upper price for producer $n$
@@ II-A Producers
Each producer $n\in\mathcal{N}$ has a capacity $\Delta^{max}_{n}\geq 0$ and a marginal price $\pi^{min}_{n}\geq 0$, and can either submit a single price (with full capacity) or a finite sequence of price and quantity pairs satisfying the following constraints involving finite upper and lower bounds:
$\pi^{min}_{n}\leq\pi_{n,k}\leq\pi^{max}_{n},\quad\forall\;k\in\{1,\dots,K_{n}\},$ (1a)
$\Delta^{min}_{n}\leq\Delta_{n,k}\leq\Delta^{max}_{n},\quad\forall\;k\in\{1,\dots,K_{n}\},$ (1b)
$\sum_{k=1}^{K_{n}}\Delta_{n,k}\leq\Delta^{max}_{n}.$ (1c)
@@ II-C Market Clearing and KKT Reformulation
    ${\sum_{n\in\mathcal{N}(z)}}\Delta_{n}^{\max}>D_{z},\;\forall z\in\mathcal{Z}$,
@@
    ${\sum_{z^{\prime}\neq z}}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\Delta_{n}^{\max}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{\max}$
    and $D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{\max}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$.
@@
Under condition (i), each zone can satisfy its own demand; thus, there exists an interior-point solution. Under condition (ii), one zone $z\in\mathcal{Z}$ may not be able to meet its own demand, i.e., $D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{\max}>0$. In this situation, we need enough demand offers left from other zones to cover the lack of supply, thus $\sum_{z^{\prime}\neq z}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\Delta_{n}^{\max}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{max}$. Since the export constraint should hold, we get that $D_{z}-\sum_{n\in\mathcal{N}(z)}\Delta_{n}^{\max}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$. ∎
@@
We now introduce a more compact notation. First we stress that, given the dependence of the feasibility set $\mathcal{X}$ on $y$, the decision variable $x_{n}$ can be expressed as an implicit function in $y$, leading to $x_{n}\equiv x_{n}(y)$. This in turn, allows us to write the objective function as an explicit function of all producers’ decision variables, i.e. $J_{n}(y_{n},y_{-n})$. Then we let $w_{n}:=(\pi_{n,k},\Delta_{n,k},x_{n,k}^{z},\lambda_{n}^{z},\mu_{n,k},\sigma_{n},\delta_{n})\in\mathcal{W}_{n}:=[\pi^{min}_{n},\pi_{n}^{max}]^{K_{n}}\times[\tilde{\Delta}_{n},\Delta_{n}^{max}]^{K_{n}}\times[0,1]^{K_{n}\cdot Z}\times[0,+\infty]^{Z}\times[0,+\infty]^{K_{n}}\times[0,+\infty]\times[0,+\infty]$. $\mathcal{W}_{n}$ is a polyhedron with dimensionality $23K_{n}+Z(K_{n}+1)+$ 2. The set of $l_{n}$ local constraints are captured by $h_{n}:\omega_{n}\rightarrow\mathbb{R}^{l_{n}}$ and that of the $r$ shared constraints by $G:\Omega\rightarrow\mathbb{R}^{r}$, where $\Omega_{n}\supseteq\mathcal{W}_{n}$ and $\Omega:=\prod_{n=1}^{N}\Omega_{n}$. In particular $h_{n}$ and $G$ are continuously differentiable mappings. Finally, we introduce the local feasibility set for each player $n$: $\widetilde{\mathcal{W}}_{n}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0\}$.
@@ III-B Nonconvex Game with Side Constraints
$0\leq\rho^{\star}\perp G(w^{\star})\leq 0$
@@ IV-A Gauss-Seidel Best Response
4:         Solve regularized Eq. (13) for agent $n$ with $[w_{1}^{k+1},\dots,\allowbreak w_{n}^{k},\dots,\allowbreak w_{N}^{k+1}]$
@@ IV-B Multi Agent Deep Reinforcement Learning
The core idea of reinforcement learning is to have an agent learning directly from exposure to an unknown environment. It is formally modeled as a Markov Decision Process (MDP) $\{S,A,r\}$ where $S$ is the state space, $A$ the action space and $r:S\times A\to\mathbb{R}$ the reward function. We define the discounted return over a trajectory of length $L$ as $R(s,a):=\sum_{t=0}^{t=L-1}\gamma^{t}r(s_{t},a_{t})$, where $10<\gamma<$ 1 is the discount factor. An agent interacts with the environment sampling action from its policy $a\sim\mu(s)$, and it aims at maximizing the reward, so the goal is to find the policy that maximizes the discounted return over an infinite horizon. In particular, defining the $Q$-function as:
@@ V Simulation Results
We now discuss the numerical results of the three algorithmic solutions we propose. We test these algorithms on real data from the German and Austrian ancillary service markets, for the period between January 2024 and August 2024 [22]. The maximum length for the bid sequence is 5, and each submitted capacity must be $\geq 5$ MW. As shown in Tab. II, the demand in Germany is much higher and volatile than that in Austria. Furthermore, the export limit is fixed at $80$ MW for both zones and the core portion is $100$ MW for Austria and 00 MW for Germany. The symmetric and low export limit suggests that producers in Germany could have some impact in Austria, while the converse situation is less likely to arise.
@@
Producer Zone $\Delta^{max}$ (MW) $\pi^{min}$
@@ V-B MARL Solution and Classical Solutions
Notice that every $\mu$ must be always 00, since every other term is multiplied by $\Delta$ and can be grouped. Furthermore, $\delta$ will be 00 for most realistic cases, since the core portion is usually a small fraction of the total demand, meaning that the constraint is always saturated. $\sigma$ on the other hand may be different from 00, which means that $\lambda$ must be different from 00 too, given the signs in the equation. Lastly, we highlight that $\pi$ is indexed by agent and bid, while $\lambda$ by zone, which means that, in most cases, the prices have to be equal to $\lambda$, virtually creating a pay-as-clear market. On the other hand, when using the Best Response algorithm, every agent computes its own $\lambda$, meaning that they are not forced to bid at the same price, and they can actually bid at very high prices with no repercussions. This constraint creates two very different markets, providing two different solutions to analyze.
@@ Appendix E MADDPG Implementation Details
  • Price Output: each price bid must be larger than the marginal price $\pi^{min}_{n}$ and smaller than the maximum bid $\pi^{max}_{n}$
@@
  • Price Module: The price module produces two values $\alpha,\beta\geq 0$, which are used to obtain the lowest and highest price of the bidding sequence as: $\pi_{0}=\min(\pi^{max},\pi^{min}+\min(\alpha,\beta))$ and $\pi_{K}=\min(\pi^{max},\pi^{min}+\max(\alpha,\beta))$. We then compute the other prices interpolating between $\pi_{0}$ and $\pi_{K}$.
@@
$r^{t}_{n}=\sum_{z\in\mathcal{Z}}\left(\frac{p_{n,z}^{t}}{d^{t}_{z}/|\mathcal{N}(z)|}-\beta\sum_{k=1}^{K_{n}}\mathbb{I}[\lambda_{z}^{t}>\pi_{n}^{min}]\cdot(\pi_{n}^{t,k}-\lambda_{z}^{t})\right)$ (20)

### v3

# Game Theory and Multi-Agent Reinforcement Learning for Zonal Ancillary Markets

@@ I-A Related Work
$\underline{\Delta}_{n},\bar{\Delta}_{n}$ Lower and upper capacity for producer $n$
@@
$\underline{\pi}_{n},\bar{\pi}_{n}$ Lower and upper price for producer $n$
@@ II-A Producers
Each producer $n\in\mathcal{N}$ has a capacity $\bar{\Delta}_{n}\geq 0$ and a marginal price $\underline{\pi}_{n}\geq 0$, and can either submit a single price (with full capacity) or a finite sequence of price and quantity pairs satisfying the following constraints involving finite upper and lower bounds:
$\underline{\pi}_{n}\leq\pi_{n,k}\leq\bar{\pi}_{n},\quad\forall\;k\in\{1,\dots,K_{n}\},$ (1a)
$\underline{\Delta}_{n}\leq\Delta_{n,k}\leq\bar{\Delta}_{n},\quad\forall\;k\in\{1,\dots,K_{n}\},$ (1b)
$\sum_{k=1}^{K_{n}}\Delta_{n,k}\leq\bar{\Delta}_{n}.$ (1c)
@@ II-C Market Clearing and KKT Reformulation
    ${\sum_{n\in\mathcal{N}(z)}}\bar{\Delta}_{n}>D_{z},\;\forall z\in\mathcal{Z}$,
@@
    ${\sum_{z^{\prime}\neq z}}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\bar{\Delta}_{n}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}$
    and $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$.
@@
Under condition (i), each zone can satisfy its own demand; thus, there exists an interior-point solution. Under condition (ii), one zone $z\in\mathcal{Z}$ may not be able to meet its own demand, i.e., $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}>0$. In this situation, we need enough demand offers left from other zones to cover the lack of supply, thus $\sum_{z^{\prime}\neq z}\left(\sum_{n\in\mathcal{N}(z^{\prime})}\bar{\Delta}_{n}-D_{z^{\prime}}\right)>D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}$. Since the export constraint should hold, we get that $D_{z}-\sum_{n\in\mathcal{N}(z)}\bar{\Delta}_{n}\leq\sum_{z^{\prime}\neq z}E_{z^{\prime}}$. ∎
@@
We now introduce a more compact notation. First we stress that, given the dependence of the feasibility set $\mathcal{X}$ on $y$, the decision variable $x_{n}$ can be expressed as an implicit function in $y$, leading to $x_{n}\equiv x_{n}(y)$. This in turn, allows us to write the objective function as an explicit function of all producers’ decision variables, i.e. $J_{n}(y_{n},y_{-n})$. Then we let $w_{n}:=(\pi_{n,k},\Delta_{n,k},x_{n,k}^{z},\lambda_{n}^{z},\mu_{n,k},\sigma_{n},\delta_{n})\in\mathcal{W}_{n}:=[\underline{\pi}_{n},\bar{\pi}_{n}]^{K_{n}}\times[\tilde{\Delta}_{n},\bar{\Delta}_{n}]^{K_{n}}\times[0,1]^{K_{n}\cdot Z}\times[0,+\infty]^{Z}\times[0,+\infty]^{K_{n}}\times[0,+\infty]\times[0,+\infty]$. $\mathcal{W}_{n}$ is a polyhedron with dimensionality $3K_{n}+Z(K_{n}+1)+2$. The set of $l_{n}$ local constraints are captured by $h_{n}:\omega_{n}\rightarrow\mathbb{R}^{l_{n}}$ and that of the $r$ shared constraints by $G:\Omega\rightarrow\mathbb{R}^{r}$, where $\Omega_{n}\supseteq\mathcal{W}_{n}$ and $\Omega:=\prod_{n=1}^{N}\Omega_{n}$. In particular $h_{n}$ and $G$ are continuously differentiable mappings. Finally, we introduce the local feasibility set for each player $n$: $\widetilde{\mathcal{W}}_{n}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0\}$.
@@ III-B Nonconvex Game with Side Constraints
$0\leq\rho^{\star}\perp G(w^{\star})\leq 0.$
@@ IV-A Gauss-Seidel Best Response
4:   Solve regularized Eq. (13) for agent $n$ with $w_{1}^{k+1},\dots,\allowbreak w_{n}^{k},\dots,\allowbreak w_{N}^{k+1}]$
@@ IV-B Multi Agent Deep Reinforcement Learning
The core idea of reinforcement learning is to have an agent learning directly from exposure to an unknown environment. It is formally modeled as a Markov Decision Process (MDP) $\{S,A,r\}$ where $S$ is the state space, $A$ the action space and $r:S\times A\to\mathbb{R}$ the reward function. We define the discounted return over a trajectory of length $L$ as $R(s,a):=\sum_{t=0}^{t=L-1}\gamma^{t}r(s_{t},a_{t})$, where $0<\gamma<1$ is the discount factor. An agent interacts with the environment sampling action from its policy $a\sim\mu(s)$, and it aims at maximizing the reward, so the goal is to find the policy that maximizes the discounted return over an infinite horizon. In particular, defining the $Q$-function as:
@@ V Simulation Results
We now discuss the numerical results of the three algorithmic solutions we propose. We test these algorithms on real data from the German and Austrian ancillary service markets, for the period between January 2024 and August 2024 [22]. The maximum length for the bid sequence is 5, and each submitted capacity must be $\geq 5$ MW. As shown in Tab. II, the demand in Germany is much higher and volatile than that in Austria. Furthermore, the export limit is fixed at $80$ MW for both zones and the core portion is $100$ MW for Austria and 0 MW for Germany. The symmetric and low export limit suggests that producers in Germany could have some impact in Austria, while the converse situation is less likely to arise.
@@
Producer Zone $\bar{\Delta}$ (MW) $\underline{\pi}$
@@ V-B MARL Solution and Classical Solutions
Notice that every $\mu$ must be always 0, since every other term is multiplied by $\Delta$ and can be grouped. Furthermore, $\delta$ will be 0 for most realistic cases, since the core portion is usually a small fraction of the total demand, meaning that the constraint is always saturated. $\sigma$ on the other hand may be different from 0, which means that $\lambda$ must be different from 0 too, given the signs in the equation. Lastly, we highlight that $\pi$ is indexed by agent and bid, while $\lambda$ by zone, which means that, in most cases, the prices have to be equal to $\lambda$, virtually creating a pay-as-clear market. On the other hand, when using the Best Response algorithm, every agent computes its own $\lambda$, meaning that they are not forced to bid at the same price, and they can actually bid at very high prices with no repercussions. This constraint creates two very different markets, providing two different solutions to analyze.
@@ Appendix E MADDPG Implementation Details
  • Price Output: each price bid must be larger than the marginal price $\underline{\pi}_{n}$ and smaller than the maximum bid $\bar{\pi}_{n}$
@@
  • Price Module: The price module produces two values $\alpha,\beta\geq 0$, which are used to obtain the lowest and highest price of the bidding sequence as: $\pi_{0}=\min(\bar{\pi},\underline{\pi}+\min(\alpha,\beta))$ and $\pi_{K}=\min(\bar{\pi},\underline{\pi}+\max(\alpha,\beta))$. We then compute the other prices interpolating between $\pi_{0}$ and $\pi_{K}$.
@@
$r^{t}_{n}=\sum_{z\in\mathcal{Z}}\left(\frac{p_{n,z}^{t}}{d^{t}_{z}/|\mathcal{N}(z)|}-\beta\sum_{k=1}^{K_{n}}\mathbb{I}[\lambda_{z}^{t}>\underline{\pi}_{n}]\cdot(\pi_{n}^{t,k}-\lambda_{z}^{t})\right)$ (20)